OpenStudy (anonymous):

What is log 3 - log 8??

OpenStudy (anonymous):

log 3 - 3 log 2=.47-3(.3)=-.43

OpenStudy (tkhunny):

log(3/8) = -0.426 -- Same thing. Use the right rules and the unique answer will not object to how you found it. log(3) = 0.477 log(8) = 0.903 0.477 - 0.903 = -0.426 -- Same thing.